OpenStudy (anonymous):

2 sqrt 27- 3 sqrt 48+ 2 sqrt 18

OpenStudy (anonymous):

Okay, you can see that in the first term, you have 2* sqrt(27) = 2* sqrt( 9 * 3 ) which equals 2* 3 * sqrt(3) = 6 sqrt(3). In the second term you have 3 sqrt(48) which is 3 * sqrt(16*3) = 3 * 4 * sqrt(3) = 12 sqrt(3). In the third term, you have 2 sqrt(18) = 2 * sqrt(9 * 2) = 2 * 3 * sqrt(2) = 6 sqrt(2) So, in total, you have 6 sqrt (3) -12sqrt(3) + 6 sqrt(2) = -6*sqrt(3) + 6*sqrt(2).

OpenStudy (anonymous):

thank u soo much i totaly understand noww !